# middlehead,lastheaddefaultContinued from previous page

### v1

@@ Abstract
Inspired by the Turing test, we present a novel methodological framework to assess the extent to which a population of machines mirrors the philosophical views of a population of humans. The framework consists of three steps: (i) instructing machines to impersonate each human in the population, reflecting their backgrounds and beliefs, (ii) administering a questionnaire covering various philosophical positions to both humans and machines, and (iii) statistically analyzing the resulting responses. We apply this methodology to the debate on scientific realism, a long-standing philosophical inquiry exploring the relationship between science and reality. By considering the outcome of a survey of over 500 human participants, including both physicists and philosophers of science, we generate their machine personas using an artificial intelligence engine based on a large-language generative model. We reveal that the philosophical views of a population of machines are, on average, similar to those endorsed by a population of humans, irrespective of whether they are physicists or philosophers of science. As compared to humans, however, machines exhibit a weaker inclination toward scientific realism and a stronger coherence in their philosophical positions. Given the observed similarities between the populations of humans and machines, this methodological framework may offer unprecedented opportunities for advancing research in experimental philosophy by replacing human participants with their machine-impersonated counterparts, possibly mitigating the efficiency and reproducibility issues that affect survey-based empirical studies.
Keywords: Scientific Realism, Experimental Philosophy, Artificial Intelligence, Turing Test
@@ 1 Machines: Like or unlike humans?
The question of whether and to what extent machines are akin to humans bears implications for a wide spectrum of disciplines, including the philosophy of mind, the nature of consciousness, intelligence and language, as well as computer science [French, 2000]. This debate traces back to the celebrated Turing test — or “imitation game” — conceived by Alan Turing in 1950, which aims to assess if machines can exhibit intelligent behavior equivalent to that of humans [Turing, 1950]. In essence, in the Turing test, a human judge ought to identify, through a text-only conversation, which of the two interlocutors is the machine and which is the human. If the judge fails to reliably distinguish the nature of the interlocutors, then the machine is deemed to have passed the test, leading to the highly controversial conclusion that it can behave like a human. A number of arguments have been leveled against the Turing test as a genuine measure of human intelligence — including the Chinese Room [Searle, 1980] and the Blockhead arguments [Block, 1981] — by demonstrating that even non-intelligent systems could potentially pass it. These criticisms have stimulated the formulation of revised versions of the test, e.g., the inverted [Watt, 1996], questioning [Damassino, 2020], and total Turing tests [Harnad, 1991].
@@
The exploration of the analogies and differences between humans and machines has experienced a new renaissance following the rapid development of Artificial Intelligence (AI) technologies [Mitchell, 2024], particularly ubiquitous large-language generative models [Naveed et al., 2024],  along with their impressive contribution to science [Krenn et al., 2022, Birhane et al., 2023] and society [Weidinger et al., 2022]. To ascertain the ability of AI models to imitate humans, several investigations have implemented Turing-like tests that diverge from the original proposal. Instead of resorting to iterative, language-based interactions with a machine, such tests evaluate the performance of AI models by directly comparing their outputs against the ground truth generated by their human counterparts, without involving the mediation of a human judge. These Turing-like tests have so far been mainly restricted to tasks that do not pose any considerable challenge to humans and that deliver an outcome that can be classified as either correct or incorrect, such as visual question answering [Yan et al., 2023], image captioning [Kasai et al., 2022] and recognition [Dodge and Karam, 2017]. However, the capability of current AI models to emulate humans in activities that involve thinking in complexity — with philosophizing being a prime example — has hitherto remained largely unexplored. Addressing this issue is particularly timely, in that the artificial intelligence encoded in large-language generative models, as quantified by the number of computational parameters, has approached the biological intelligence encoded in homo sapiens, as quantified by the number of synapses in the brain [Schwartz, 2022]. Philosophizing is generally regarded as a uniquely human pursuit, embedded in a collective and multifaceted endeavor, demanding internal consistency rather than right-or-wrong classifications, and often shaped by factors that are typically inaccessible to machines. This latter aspect is corroborated by recent empirical studies indicating that philosophical judgments are influenced by, e.g.,  birth cohort [Hannikainen et al., 2018], gender [Buckwalter and Stich, 2013], alcohol consumption [Duke and Bègue, 2015], wording and ordering effects [Petrinovich and O’Neill, 1996], as well as personality traits [Bartels and Pizarro, 2011], as hypothesized by William James the beginning of the twentieth century [James, 2000].
@@
Here, we design and implement a novel methodological framework to quantify the degree to which the philosophical positions held by machines mirror those of humans. First, we devise a general protocol to instruct large-language, generative AI models to impersonate a population of humans, reflecting the diverse backgrounds and beliefs of individuals. Second, we apply this protocol to compare the philosophical positions endorsed by a population of humans with those endorsed by the corresponding population of machines. As a case study, we consider the philosophical debate on scientific realism [Psillos, 1991] — a century-long inquiry into the relationship between science and reality — owing to its central role in the philosophy of science along with its interdisciplinary relevance attested by the engagement of philosophers and scientists alike.  Our findings indicate that a population of machines holds beliefs that are, on average, quite similar to those held by the population of humans, differing only by a few percent. As compared to humans, however, machines exhibit a slightly less pronounced inclination toward scientific realism and significantly more coherent philosophical views. We additionally identify common patterns underlying human and machine populations when confronted with the philosophical challenges raised by the realism debate. Overall, our analysis unveils the ability of machines to imitate humans when addressing complex issues, possibly paving the way for the introduction of AI-assisted approaches in the rapidly emerging field of experimental philosophy.
2 Overview of selected positions in the scientific realism debate
@@
Our work builds on the approach of Henne and coworkers [Henne et al., 2024], which employs a questionnaire to survey the views of physicists and philosophers of science within the scientific realism debate. The questionnaire consists of 30 statements, as listed in Table LABEL:Table-1, describing, either directly or indirectly through specific examples, four prevailing philosophical positions, that is, scientific realism (including several forms of selective realism), instrumentalism, pluralism, and perspectivism. We briefly outline these positions.
@@
Scientific realism [Psillos, 1991] is the view that science portrays a faithful representation of the world (S1 and S2) by discovering objects that are beyond our perception (S4 and S6), such as electrons (S13, S15, S17, S18), and formulating theories that are at least approximately true (S8), such as the Big Bang theory (S20) or general relativity (S21). This view is rooted in three commitments. First, a metaphysical commitment, holding that there exists a mind-independent reality, namely, a reality that is not sensitive to our specific theories (S14) or our particular approach to manipulate and describe it (S11). Second, a semantic commitment, holding that successful scientific theories ought to be interpreted as (approximately) true by correspondence (S21), thus revealing the actual features of reality (S23), e.g., the nature of space and time (S22). Third, an epistemic commitment, holding that the intrinsic scope of science is to achieve such truth by correspondence (S10, S26).
@@
Selective realism [Chakravartty, 2010] prescribes that the ontological commitment should not be endowed to scientific theories en bloc. Instead, belief should be restricted to a narrow subset of theoretical claims. Depending on the subset of theoretical claims regarded as epistemically secure, two major yet opposed forms of selective realism have been developed, that is, structural realism and entity realism. Structural realism [Worrall, 1989, Ladyman, 1998] warrants belief in relations, as typically subsumed in the mathematical structures of scientific theories, while advocating for skepticism about entities (S27). Entity realism [Hacking, 1983, Cartwright, 1983] warrants belief in entities, especially those that are amenable to experimental manipulation, while advocating for skepticism about the mathematical structures that purport to describe them (S13, S14, S18, S19).
@@
Antithetical to scientific realism is instrumentalism [Rowbottom, 2019, Stanford, 2010], which denies that successful scientific theories can offer access to the ultimate nature of reality (S3). Rather, it acknowledges the effectiveness of scientific theories as devices to classify and predict phenomena (S9, S23), while considering their posited unobservable entities being mere fictions (S5) or constructions (S7). Instrumentalism is often driven by a ‘pessimistic meta-induction’ [Laudan, 1981] which draws from the history of science to infer that present-day theories, analogous to the superseded theories of the past, are likewise poised to be abandoned (S12). Unlike scientific realism, instrumentalism rejects the notion that scientific theories are conducive to truth by correspondence, regarding the unobservable entities existing only within the sole province of the theories that postulate them, leading to a sort of internal realism (S16).
@@
Besides the realism-instrumentalism dichotomy, the debate has given rise to additional positions, most notably pluralism and perspectivism. Pluralism [Chang, 2012, Massimi, 2018] has been articulated in two versions. Epistemic pluralism maintains that multiple and even incompatible theories can nevertheless be of equal epistemic value (S24, S25), in opposition to scientific realism which contends that the nature of the world is unique. Methodological pluralism maintains that conflicting theories are valuable for the progress of scientific inquiry (S28). Perspectivism [Massimi, 2022, Giere, 2006, Lipton, 2007] advocates that different theories entail different perspectives on the world, each of them delivering a representation from a particular and limited point of view (S30) that is influenced by the cultural traditions and historical periods in which the theories are formulated (S29).
@@
[ caption = List of statements assessed by physicists, philosophers of science, and their corresponding machine-generated personas, covering various philosophical positions that emerged in the scientific realism debate. Statements highlighted in blue (red) denote agreement with realism (instrumentalism), whereas statements highlighted in grey are compatible with both views and include pluralism and perspectivism. Adapted from the survey of Henne and coworkers [Henne et al., 2024]. , label = Table-1, ] colspec=|X[1]|X[18]|X[8]|, rowhead = 1, hlines, S & Statement Philosophical position
@@
3 Assessing the views of a population of machines
Of the 30 statements listed in Table LABEL:Table-1, 14 reflect an inclination toward scientific realism (S1, S2, S4, S6, S8, S10, S11, S13, S14, S15, S17, S18, S20, S22), 8 reflect an inclination toward instrumentalism (S3, S5, S7, S9, S12, S16, S23, S25), and the remaining 8 are compatible with both positions (S19, S20, S24, S26, S27, S28, S29, S30). To assess which views, among those presented in Table LABEL:Table-1, are endorsed by physicists and philosophers of science, Henne and coworkers [Henne et al., 2024] administered these statements to 535 participants — consisting of 384 physicists and 151 philosophers of science — requesting them to rate their agreement by assigning an ‘agreement score’ ranging from 0% (“This sounds completely wrong to me”) to 100% (“This strikes me as exactly right”). Because participants may not be familiar with the specialized terminology involved in the philosophical debate, they were instructed to assign the agreement score according to their immediate understanding of the statement (“Many of the statements may seem unclear. For example, terms like ‘truth’ and ‘reality’ can be understood in many ways. Please answer according to your immediate inclination”).

### v2

@@ Abstract
Inspired by the Turing test, we present a novel methodological framework to assess the extent to which a population of machines mirrors the philosophical views of a population of humans. The framework consists of three steps: (i) instructing machines to impersonate each human in the population, reflecting their backgrounds and beliefs, (ii) administering a questionnaire covering various philosophical positions to both humans and machines, and (iii) statistically analyzing the resulting responses. We apply this methodology to the debate on scientific realism, a long-standing philosophical inquiry exploring the relationship between science and reality. By considering the outcome of a survey of over 500 human participants, including both physicists and philosophers of science, we generate their machine personas using an artificial intelligence engine based on a large language model. We reveal that the philosophical views of a population of machines are, on average, similar to those endorsed by a population of humans, irrespective of whether they are physicists or philosophers of science. As compared to humans, however, machines exhibit a weaker inclination toward scientific realism and a stronger coherence in their philosophical positions. Given the observed similarities between the populations of humans and machines, this methodological framework may offer unprecedented opportunities for advancing research in the empirical social sciences by complementing human participants with their machine-impersonated counterparts.
Keywords: Artificial Intelligence, Turing Test, Scientific Realism, Philosophy, Physics.
@@ 1 Machines: Like or unlike humans?
The question of whether and to what extent machines are akin to humans bears implications for a wide spectrum of disciplines, including the philosophy of mind, the nature of consciousness, intelligence and language, as well as computer science [French, 2000]. The canonical beginning of the discussion is credited to Alan Turing who, in his now-classic 1950 essay, posed the question “Can machines think?” [Turing, 1950]. Turing argues that, formulated as such, the question is “too meaningless” and thus advocates for its replacement with an actionable protocol “which is closely related to it:” the Turing test or imitation game. In the Turing test, a human judge ought to identify, through a text-only conversation, which of the two interlocutors is the machine and which is the human. If the judge fails to reliably distinguish the nature of the interlocutors, then the machine is said to have passed the test, leading to the highly controversial conclusion that it can exhibit an intelligent or human-like behavior. A number of arguments have been leveled against the Turing test as a genuine measure of human intelligence, such as the Chinese Room [Searle, 1980] and the Blockhead arguments [Block, 1981], claiming that even non-intelligent systems could potentially pass it. This suggests that inner psychology cannot be merely reduced to the observation of the outward behavior of an agent. Some criticisms have further stimulated the formulation of revised versions of the test, e.g., the inverted [Watt, 1996], questioning [Damassino, 2020], and total Turing tests [Harnad, 1991].
The development of approaches to detect the presence of thought in putatively minded agents traces back to Cartesian philosophy. In the Discourse on the Method, Descartes proposes an approach for distinguishing humans from automata, arguing that the latter are invariably incapable of being convincingly disguised as the former. Gunderson has noted that Descartes in fact distinguishes two tests [Gunderson, 1964]. First, the language test, whereby a machine “could never use speech or other signs placing our thoughts on record for the benefit of others,” for “it never happens that it arranges its speech in various way, in order to reply appropriately to everything that may be said in its presence, as even the lowest type of man can do.” Second, the action test, which is based on the idea that “although machines can perform certain things as well as or perhaps better than any of us can do, they infallibly fall short in others, by the which means we may discover that they did not act from knowledge, but only from the disposition of their organs. For while reason is a universal instrument which can serve for all contingencies, these organs have need of some special adaptation for every particular action. From this it follows that it is morally impossible that there should be sufficient diversity in any machine to allow it to act in all events of life in the same way as our reason causes us to act” [Gunderson, 1964]. Erion has offered an alternative interpretation of the action test by arguing that it has to be regarded as a test of common sense, understood as the ability to perform tasks that even the most simple-minded adult human can do [Erion, 2001]. Be that as it may, Turing was aware of the Cartesian language test, which played an important role in the introduction of the imitation game [Abramson, 2011].
@@
The exploration of the analogies and differences between humans and machines has recently experienced a new renaissance following the rapid progress of generative Artificial Intelligence (AI) systems [Mitchell, 2024] such as large language models [Naveed et al., 2024]—a class of neural networks trained on vast amounts of text data to understand, analyze, and generate natural language—along with their impressive contribution to science [Krenn et al., 2022, Birhane et al., 2023] and society [Weidinger et al., 2022]. Large language models appear to instantiate what Descartes in the Discourse deemed “not conceivable,” namely, the realization of “a machine” that “should produce different arrangements of words so as to give an appropriately meaningful answer to whatever is said in its presence, as the dullest of men can do.” To ascertain the ability of AI systems to imitate humans, several investigations have implemented Turing-like tests that diverge from the original proposal. Instead of resorting to iterative, language-based interactions with a machine, such tests evaluate the performance of AI models by directly comparing their outputs against the ground truth [^1] Here, ‘ground truth’ is used in the statistical sense—not the philosophical one—referring to the ideal expected result against which the performance of a machine is evaluated. generated by their human counterparts, without involving the mediation of a human judge. These Turing-like tests have so far been mainly restricted to tasks that do not pose any considerable challenge to humans and that deliver an outcome that can be classified as either correct or incorrect, such as visual question answering [Yan et al., 2023], image captioning [Kasai et al., 2022] and recognition [Dodge and Karam, 2017]. However, the capability of current AI models to emulate humans in activities that involve thinking in complexity has hitherto remained largely unexplored. Addressing this issue is particularly timely, in that the artificial intelligence encoded in large language generative models, as quantified by the number of computational parameters, has approached the biological intelligence encoded in homo sapiens, as quantified by the number of synapses in the brain [Schwartz, 2022].
Here, we consider the question: “Can machines philosophize?” Philosophizing is generally regarded as a uniquely human pursuit, embedded in a collective and multifaceted endeavor, demanding, for example, internal consistency rather than right-or-wrong classifications, and often shaped by factors that are typically inaccessible to machines. This latter aspect is corroborated by recent empirical studies indicating that philosophical judgments are influenced by, e.g.,  birth cohort [Hannikainen et al., 2018], gender [Buckwalter and Stich, 2013], alcohol consumption [Duke and Bègue, 2015], wording and ordering effects [Petrinovich and O’Neill, 1996], as well as personality traits [Bartels and Pizarro, 2011], as hypothesized by William James the beginning of the twentieth century [James, 2000]. [^2] We do not take a normative stance on whether such influences should or should not occur, as this question is irrelevant the scope of the present work. Inspired by the Turing test, we design and implement a novel methodological framework to quantify the degree to which the philosophical positions held by machines mirror those of humans.
The rationale of our study is twofold, encompassing both fundamental and applied dimensions. On the fundamental side, previous works comparing humans and machines—typically framed in the tradition of the Turing test—have largely concentrated on the assessment of individual agents. By contrast, our approach shifts the focus from individuals to populations, asking whether AI can emulate the statistical distribution of views that emerges in human groups. This population-level perspective not only extends the methodological scope of the imitation game but also opens novel questions about the capacity of machines to reproduce patterns of collective human reasoning. On the applied side, our framework may offer practical benefits for quantitative research in the empirical social sciences, where survey-based methods are widely used to chart attitudes and inform debates. In these contexts, machine-generated populations with controlled characteristics could complement, or in some cases partially substitute for, human subject pools, thereby broadening the methodological toolkit available to researchers while helping overcome some challenges that are inherent to survey-based studies, such as recruiting a sufficiently large and representative sample, high costs and long timelines required to gather responses, or survey design, given that poorly worded questions may only become evident once responses are collected.
In this work, first we devise a general protocol to instruct large language, generative AI models to impersonate a population of humans, reflecting the diverse backgrounds and beliefs of individuals. Second, we apply this protocol to compare the philosophical positions endorsed by a population of humans with those endorsed by the corresponding population of machines. As a case study, we consider the debate on scientific realism [Psillos, 1999]—a century-long inquiry into the relationship between science and reality—owing to its central role in the philosophy of science.  This choice is guided by both philosophical and empirical considerations. From a philosophical perspective, scientific realism stands as one of the most enduring and contested issues in philosophy, engaging the sustained attention of both scientists and philosophers. Its centrality and unresolved status make it an especially informative case for evaluating whether machines are capable of navigating and reproducing reasoning within a live, interdisciplinary controversy. Indeed, unlike many applications where AI outputs are checked against an objective ground truth (e.g., image captioning), issues that emerged within the scientific realism debate typically involve multiple, coexisting perspectives rather than a single, definitive answer. From an empirical perspective, our study builds upon the work of Henne and coworkers [Henne et al., 2024], which offers one of the most recent and comprehensive surveys available on attitudes toward scientific realism. This survey not only provides a robust benchmark for evaluating AI-generated responses but also encompasses a heterogeneous population, including physicists from a range of subfields and philosophers subscribing to diverse positions. Such diversity allows us to examine whether machines can approximate not only the aggregate tendencies of human respondents but also the more fine-grained perspectives of different disciplinary groups.
@@
Our findings indicate that a population of machines holds beliefs that are, on average, quite similar to those held by the population of humans, differing only by a few percent. As compared to humans, however, machines exhibit a slightly less pronounced inclination toward scientific realism and significantly more coherent philosophical views. We additionally identify common patterns underlying human and machine populations when confronted with the philosophical challenges raised by the realism debate. Overall, our analysis unveils the ability of machines to imitate human populations when addressing complex issues, possibly paving the way for the introduction of AI-assisted approaches in the empirical social sciences.
2 Philosophical views in the scientific realism debate
@@
Our work builds on the approach of Henne and coworkers [Henne et al., 2024], which employs a questionnaire to survey the views of physicists and philosophers of science within the scientific realism debate. The questionnaire consists of 30 statements listed in Table LABEL:Table-1 describing, either directly or indirectly through specific examples, four philosophical positions, that is, scientific realism (including several forms of selective realism), instrumentalism, pluralism, and perspectivism. We briefly outline these positions.
@@
Scientific realism [Psillos, 1999] is the view that science portrays a faithful representation of the world (S1 and S2) by discovering objects that are beyond our perception (S4 and S6), such as electrons (S13, S15, S17, S18), and formulating theories that are at least approximately true (S8), such as the Big Bang theory, as an example of speculative physics (S20), [^3] Speculative physics involves theoretical frameworks and conjectures to explore phenomena beyond currently available empirical evidence and testability. or general relativity (S21). This view is rooted in three commitments. First, a metaphysical commitment, holding that there exists a mind-independent reality, namely, a reality that is not sensitive to our specific theories (S14) or our particular approach to manipulate and describe it (S11). Second, a semantic commitment, holding that successful scientific theories ought to be interpreted as (approximately) true by correspondence (S21), thus revealing the actual features of reality (S23), e.g., the nature of space and time (S22). Third, an epistemic commitment, holding that belief in scientific theories is justified by compelling epistemic reasons. These commitments are often complemented by the axiological claim that the scope of science is to achieve truth by correspondence (S10).
@@
A prevailing variety of scientific realism is selective realism [Chakravartty, 2010], which prescribes that that the ontological commitment should not be endowed to scientific theories en bloc. Instead, belief should be restricted to a narrow subset of theoretical claims. Depending on the subset of theoretical claims regarded as epistemically secure, three main forms of selective realism have been developed, that is, deployment realism [Psillos, 1999], structural realism, and entity realism. Structural realism [Worrall, 1989, Ladyman, 1998] warrants belief in relations, as typically subsumed in the mathematical structures of scientific theories, while advocating for skepticism about entities (S27). Entity realism [Hacking, 1983, Cartwright, 1983] warrants belief in entities, especially those that are amenable to experimental manipulation, while advocating for skepticism about the mathematical structures that purport to describe them (S13, S14, S18, S19).
@@
One of the views opposing scientific realism is instrumentalism [Rowbottom, 2019, Stanford, 2010], [^4] Other relevant antirealist views are, e.g., positivism and, more recently, constructive empiricism [van Fraassen, 1980] which denies that successful scientific theories can offer access to the ultimate nature of reality (S3). Rather, it acknowledges the effectiveness of scientific theories as devices to classify and predict phenomena (S9, S23), while considering their posited unobservable entities being mere fictions (S5) or constructions (S7). Instrumentalism is often driven by a ‘pessimistic meta-induction’ [Laudan, 1981] which draws from the history of science to infer that present-day theories, analogous to the superseded theories of the past, are likewise poised to be abandoned (S12). Unlike scientific realism, instrumentalism rejects the notion that scientific theories are conducive to truth by correspondence, regarding the unobservable entities existing only within the sole province of the theories that postulate them.
@@
Besides the realism-instrumentalism dichotomy, the debate has given rise to additional positions, such as pluralism and perspectivism. Pluralism [Chang, 2012, Massimi, 2018] has been articulated in many flavors, some compatible with some forms of realism, other not. Epistemic pluralism, for instance, maintains that multiple and even incompatible theories can nevertheless be of equal epistemic value, while ontological pluralism embraces the idea that the nature of the world is not unique (S24, S25). Methodological pluralism maintains that conflicting theories are valuable for the progress of scientific inquiry (S28). Perspectivism [Massimi, 2022, Giere, 2006, Lipton, 2007] advocates that different theories entail different perspectives on the world, each of them delivering a representation from a particular and limited point of view (S30) that is influenced by the cultural traditions and historical periods in which the theories are formulated (S29). Closely related to these position is internal realism [Putnam, 1981] which rejects the ‘God’s-Eye Point of View’ inherent in metaphysical realism—the view that the external reality is objective and independent of how inquiring agents conceptualizes it—holding instead that our understanding and the world and the truth of our theories are relative to a given conceptual scheme (S16).
@@
[ caption = List of statements assessed by physicists, philosophers of science, and their corresponding machine-generated personas, covering various philosophical positions that emerged in the scientific realism debate. Statements highlighted in blue (red) denote agreement (disagreement) with scientific realism, whereas statements highlighted in grey may be compatible with both views and include pluralism and perspectivism. Reproduced verbatim from the survey of Henne and coworkers [Henne et al., 2024], label = Table-1, ] colspec=|X[1]|X[18]|X[8]|, rowhead = 1, hlines, S Statement Philosophical position
@@
3 Assessing the views of machines
Of the 30 statements listed in Table LABEL:Table-1, 14 reflect an inclination toward scientific realism (S1, S2, S4, S6, S8, S10, S11, S13, S14, S15, S17, S18, S20, S22), 8 reflect an opposition to realism (S3, S5, S7, S9, S12, S16, S23, S25), while the remaining 8 are compatible with both positions (S19, S20, S24, S26, S27, S28, S29, S30). To assess which views, among those presented in Table LABEL:Table-1, are endorsed by physicists and philosophers of science, Henne and coworkers [Henne et al., 2024] administered these statements to 535 participants—consisting of 384 physicists and 151 philosophers of science—requesting them to rate their agreement by assigning an ‘agreement score’ ranging from 0% (“This sounds completely wrong to me”) to 100% (“This strikes me as exactly right”). Because participants may not be familiar with the specialized terminology involved in the philosophical debate, they were instructed to assign the agreement score according to their immediate understanding of the statement (“Many of the statements may seem unclear. For example, terms like ‘truth’ and ‘reality’ can be understood in many ways. Please answer according to your immediate inclination”).